### v2

# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

Abstract
We study the Compton-rocket effect of strong radiation force accelerating electrons in an opaque fireshell (or fire spot) of dense photons and electron-positron pairs, whose temperature is spatially inhomogeneous and exceeds the electron mass. We find the possibility of the charged-particle acceleration and the avalanche runaway process, leading to a non-trivial probability of ultra-high-energy (UHE) electrons and protons, which subsequently produce very-high-energy (VHE) photons and neutrinos. In a simplified one-dimensional model, we qualitatively show such peculiar dynamics using the fireball, Gamma-Ray Burst central engine, whose inner part inflows and forms a gravitationally trapped fireshell (halo) around the horizon of a black hole. The fireshell is metastable, cooling via UHE particle emissions and blackbody radiation. We calculate the UHE particle luminosity varying in time, and discuss the peculiar features of such produced UHE particles, which lead to VHE particles, in connection with possible numerical simulations, observations and experiments.
1 Introduction
The discovery of ultra-high-energy (UHE $\sim 10^{18-20}$ eV) charged particles (cosmic rays) has been for decades. The acceleration mechanism of UHE charged particles remains mysterious, although much theoretical and observational progress has been made to understand their nature and identify their sources (see review [1]). Very-high-energy (VHE $\sim 10^{12-15}$ eV) neutral particles (photons and neutrinos) have been recently observed by the LHAASO [2], MAGIC [3], and IceCube [4] collaborations. Much theoretical attention has been given to understanding the physical origin of UHE charged particles and VHE neutral particles, and identifying astrophysical sources where they come from.
It is physically reasonable to assume that these highly energetic particles may originate from compact and powerful astrophysical sources, such as accretion disks, massive core collapses, binary coalescence and active galactic nuclei. In these sources, there may exist highly energetic arenas, where strong gravitational fields, dense radiation fields, and charged particles interact non-linearly with one another. These interactions possibly result in the high-energy processes of accelerating charged particles, leading to UHE and VHE particles.
One of these powerful astrophysical sources can be the central engine for Gamma-ray bursts (GRBs), an opaque fireballs of large ultra-relativistic particle (photon and pair) energy density and pressure [5], and see reviews [6, 7, 8, 9, 10, 11, 12, 13]. Within the general relativity (GR) framework, we studied [14, 15] the hydrodynamical outflow of fireballs by using also the Boltzmann rate equation for electrons and positrons (pairs) annihilation to photons, obtaining the relation between the photon heat and baryon kinetic energies for main bursts and afterglow radiations. We will first use the fireball, Gamma-Ray Bursts (GRBs) central engine, to show a possible highly energetic arena, a fireshell or fire spot, then illustrate our studies and results of the high-energy interactions and processes that produce UHE and VHE particles in such an arena.
We will study in Section 2 that through the gravitational force and thermal pressure, the fireball’s inner part (halo around a black hole) is trapped and forms a metastable thermal fluid of photons and pairs at high energy density and temperature. Using such an opaque horizon halo, we study the Compton-rocket effect on accelerating charged particles (Section 3) and the avalanche runaway process (Section 4) leading to UHE electrons and protons. We calculate the luminosity of UHE particles and the halo cooling in Section 6, and discuss the peculiar features of UHE and VHE particles produced in this dynamics, in connection with numerical simulations, observations, and experiments in the final Section 7. We use the $\hbar=c=1$ and Compton unit of the electron mass $m_{e}$, unless otherwise specified in necessary cases for clarification.
Figure 1: Figure 2 of Ref. [16] shows the world lines of fireball hydrodynamics expanding outflow (dot line) and trapped inflow (solid line) in the Schwarzschild geometry of a collapsing massive stellar core (dashed line), which forms a black hole of mass $M$. Near the horizon radius $r_{+}=2M$ of a formed black hole, the gravitational time dilation $\delta t\gg\delta\tau$ for an observer at a distance is evident; $\delta\tau$ is a finite proper time (distance) for a local observer. The separatrix radius between out and in flows is $\bar{R}=4M$, and the trapped inflow energy is comparable with the expanding outflow energy. The total inner and outer fireball energy gain from the collapsing system’s gravitational binding energy, and it should be much smaller than $(M^{2}/r_{\infty}-M^{2}/r_{+})/2=M/4$ in the Newtonian approximation.
2 Trapped fireshell (halo) around black hole horizon
2.1 Trapped fireshell (halo) of energetic and dense photons and pairs
Although the dynamics of fireball formation are still under investigation  [^1] In Refs. [17, 18], we present the detailed studies of the possibility of how fireballs or jets form during gravitational collapses of massive rotating stellar cores or binary coalescence: heavy hadrons gain kinetic energy from gravitational potential, and their heavy-ion collisions produce photons and pairs, resulting in the fireball or jet formation., the hydrodynamical outflow of fireballs accounting for GRBs has been well understood upon the assumptions that entire fireballs are propelled out in a flat spacetime. Nevertheless, a scrutiny GR study [16] investigates the hydrodynamical evolution of fireballs in the Schwarzschild geometry of a massive stellar core, which undergoes a gravitational collapse process to form a black hole of mass $M$. The result shows in Fig. 1 that the fireball’s inner part $(r<\bar{R})$ is trapped due to strong gravitational attraction, prevailing over thermal pressure repulsion. On the contrary, its outer part $(r>\bar{R})$, where gravitational attraction is weak and thermal pressure repulsion prevails, hydrodynamically outflows, undergoing internal and external shocks in its evolution, accounting for the GRB phenomenon as usually described. The radius separatix between the inner and outer parts is $\bar{R}\approx 4GM=2r_{+}$ near the black hole horizon $r_{+}=2M$  [^2] We can approximately estimate the separatrix $\bar{R}$ based on the balance between thermal repulsive pressure and gravitational attractive force $4\pi r_{+}^{2}\rho_{\gamma}(r_{+})c\sim Gm_{\gamma}M/r_{+}^{2}$ on the fireball inner part of energy mass $m_{\gamma}\sim\rho_{\gamma}(r_{+})4\pi r_{+}^{2}\bar{R}$ and fireball density $\rho_{\gamma}(r_{+})$ at the horizon..
The gravitational collapsing system comprises a cold baryon core and a hot fireball of relativistic particles. When the spacetime horizon is approaching, the final phase of the gravitational collapse of the nonrelativistic baryon matter core and the relativistic particles’ fluid (fireball inner part), which becomes denser, must be nontrivial. It can neither be a free-fall process of the time scale $\sim GM/c$, nor a rapid transit process, like a GRB prompt emission of an expanding fireball at its transparency. In the final phase of gravitational collapses, the gravitational time dilation effect is significant for an observer at a distance, see Fig. 1 and its caption.
The trapped fireball’s inner part, which we call the fireshell (or halo) for short, is highly energetic and dense, comprising numerous ultra-relativistic particles and antiparticles, which are represented by photons $\gamma$, electrons $e^{-}$ and positrons $e^{+}$ pairs. They have zero chemical potential and are electrically neutral in total. It is completely opaque, undergoing the rapid back-and-forth process
$\gamma+\gamma\leftrightarrow e^{+}+e^{-},$ (2.1)
which establishes a local thermal equilibrium at high temperatures and densities. Since the spacetime length scales of the back-and-forth process (2.1) are very small, we describe all photons $\gamma$ and pairs $e^{+}e^{-}$ together as a radiation fluid, rather than identifying them as individual photons, electrons $e^{-}$ and positrons $e^{+}$  [^3] In other words, relativistic fermions always pair with their antiparticles via exchanging two photons. Therefore, we ignore the notion of anti-particles like positrons $e^{+}$ in this article.. We treat a local element of the trapped fireball as a perfect radiation fluid of large energy density $\rho_{\gamma}(r)$ and pressure $P_{\gamma}(r)$ with the Equation of State $P_{\gamma}(r)=\rho_{\gamma}(r)/3$. They are inhomogeneous as functions of radius $r$, due to the presence of the gravitational field of a massive collapsing stellar core or black hole. Once more, we stress that the total thermal energy of the trapped fireball’s inner part is comparable to that of the expanding fireball’s outer part for GRBs [16].
Moreover, in distinct contrast with the gravitational collapse processes of pressure-less baryon dust, the dynamics of inflow contraction of the opaque fireshell is much more complex, because the gravitational potential, the fireshell density and pressure increase. To study such an inflow contraction in time, we should, in principle, consider the inhomogeneous radiative fluid and Einstein equation in the spherically symmetric Lemaître spacetime [19]  [^4] The Lemaître metric reduces to the Lemaître-Tolman-Bondi one in the pressureless dust case, whose solutions are well known. However, despite many years of work, the solution of the Lemaître metric with a general perfect fluid source is not known. in the presence of a black hole, whose metric elements are functions of the coordinate radius $r$ and time $t$. Observe that (i) the fireshell inflow contraction should be slow in time because of its pressure being very large; (ii) the fireshell is deeply opaque so that its thermal energy dissipation rate is small. Therefore, the fireshell in Lemaître spacetime should be a metastable configuration undergoing slow contraction adiabatically. Namely, the Lemaître metric elements, the fireshell energy density and pressure should be slowly varying in time. Thus, we approximately adopt the static limit, all quantities being independent of time, the Lemaître spacetime reduces to the Tolman–Oppenheimer–Volkoff (TOV) one [20] in the interior of the fireshell with the exterior in Schwarzschild coordinates.
2.2 Tolman–Oppenheimer–Volkoff equation for halo energy density
In this section, we adopt the spherically symmetric TOV equation for the fireshell to study the gravitational equilibrium configuration of ultra-relativistic photons and pairs trapped in the Schwarzschild spacetime near the horizon,
${\frac{dP_{\gamma}}{dr}}=-{\frac{Gm(r)}{r^{2}}}\rho_{\gamma}\left(1+{\frac{P_{\gamma}}{\rho_{\gamma}}}\right)\left(1+\frac{4\pi r^{3}P_{\gamma}}{m(r)}\right)\left(1-{\frac{2Gm(r)}{r}}\right)^{-1},$ (2.2)
in the Schwarzschild geometry of radial coordinate $r$ and time $t$ of an observer at infinity. The TOV equation describes the equilibrium balance between the gravitational and radiation forces. Within the sphere of a radius $r$, the total energy mass ${\textstyle m(r)=M+E_{\gamma}(r)}\approx M$ and the photon-pair fluid energy $E_{\gamma}(r)$ is much smaller than the black hole mass $M$.
Figure 2: Left: the trapped fireshell (halo) energy density $\rho_{\gamma}$, solution to (2.2), is expressed in units of $m_{e}^{4}$ and as functions of radius $r$, which is in the unit of the horizon radius $r_{+}=2GM=2.33\times 10^{6}$ cm for the black hole mass $M=7.75M_{\odot}$ (2.4). Right: the trapped fireshell (halo) temperature $T_{\gamma}=(\rho_{\gamma}/\sigma_{{}_{\rm SB}}d_{f})^{1/4}$ (2.9) is expressed in units of the electron mass $m_{e}\approx 0.5{\rm MeV}$. The maximal temperature $T^{+}_{\gamma}$ near the horizon rapidly decreases to the asymptotic value $\approx 2m_{e}$ near the boundary $2r_{+}=4GM$, where the electron-positron pair annihilation takes place. The Compton energy density $m_{e}^{4}\approx 4.54\times 10^{21}{\rm erg/cm}^{3}$ is of the same order as the critical energy density $\rho_{c}=E_{c}^{2}/(8\pi)=5.45m_{e}^{4}$, where $E_{c}=m_{e}^{2}/e$ is the critical electric field for electron and positron pair production in the semi-classical approximation [12].
Rescaling the radius $r$ in the unit of $2GM$ and density $\rho_{\gamma}$ in the unit of “average density” $\rho_{{}_{M}}=M/V_{M}$, where “volume” $V_{M}=(4\pi/3)(2GM)^{3}$, we recast the TOV equation (2.2) as a dimensionless equation
${\frac{d}{dr}}\left(\frac{\rho_{\gamma}}{\rho_{{}_{M}}}\right)=-\frac{3}{2r^{2}}\left(\frac{\rho_{\gamma}}{\rho_{{}_{M}}}\right)\left[1+r^{3}\left(\frac{\rho_{\gamma}}{\rho_{{}_{M}}}\right)\right]\left(1-{\frac{1}{r}}\right)^{-1},$ (2.3)
for the density $\rho_{\gamma}$. We consider a massive stellar baryon core of neutron stars at the nuclear saturation density $\rho_{n}\approx 0.16\,{\rm fm}^{-3}=1.6\times 10^{38}/{\rm cm}^{3}$, gravitationally collapsing to a compact object of size about $2GM$ of the black hole horizon. The “average density” $\rho_{{}_{M}}=M/V_{M}$ varies from $(2-10)\rho_{n}$ for the mass $M$ varying from $(3-10)M_{\odot}$. The value $\rho_{{}_{M}}=10\rho_{n}$ effectively corresponds to a black hole of the mass $M$ and horizon radius $r_{+}$
$M=7.75M_{\odot},\quad r_{+}=2GM=2.33\times 10^{6}{\rm cm}.$ (2.4)
The photon-pair pressure $P_{\gamma}(r)=\rho_{\gamma}(r)/3$ should be maximum near the horizon, where the gravitational force is the strongest. To illustrate the fireshell properties, we choose the photon-pair density near the horizon
$\rho_{\gamma}^{+}\equiv\rho_{\gamma}(r_{+})=\eta\rho_{{}_{M}}=\rho_{n},~~~\eta=0.1,$ (2.5)
as the boundary conditions for the TOV equation (2.3). The reason for $\rho_{\gamma}^{+}\sim\rho_{n}$ is that at high energies, many species of relativistic particles, lepton-antilepton pairs and photons, light quark-antiquark pairs and gluons, could be excited, contributing to the energy density $\rho_{\gamma}^{+}$. It turns out to be a fluid of lepton(pair)-photon and quark(pair)-gluon plasma. For the sake of simplified notations, we use the photon ($\gamma$) and pair ($e^{-}e^{+}$) notations representing all possible gauge bosons and fermion-antifermion pairs, and use the subscript $\gamma$” for the quantities of the trapped fireshell of photons and pairs.
Solving the TOV equation (2.3) with boundary conditions (2.5), we show in Fig. 2 (left) that the energy density $\rho_{\gamma}$ of photons and pairs is maximal near the horizon $r_{+}$ and rapidly decreases as the radius increases to $2r_{+}$. The “trapped fireshell (halo)” is named for such an energetic radiation fluid configuration of photons and pairs near the black hole horizon, where the maximal gravity effect confines the opaque radiation fluid of photons and pairs at high densities. The macroscopic time scale of hydrodynamical relaxation for establishing a trapped fireshell is about $\bar{R}/v_{s}$, where the fluid sound velocity $v_{s}=(P_{\gamma}/\rho_{\gamma})^{1/2}=c/\sqrt{3}$.
The total trapped fireshell energy is an integration over the energy density $\rho_{\gamma}$ from the horizon $r_{+}=2GM$ up to the boundary $\bar{R}=4GM=2r_{+}$
$E_{\gamma}$ $=$ $\int_{r_{+}}^{2r_{+}}\rho_{\gamma}(r)\frac{4\pi r^{2}dr}{(1-\frac{r_{+}}{r})^{1/2}}=\rho_{\gamma}^{+}4\pi r_{+}^{2}\delta\ell,$ (2.6)
which is the thermal energy of photons and pairs, as will be discussed in the next section. Note that the trapped thermal energy is not small, but comparable with the expanding outflow energy for GRBs, see Fig. 2 caption and Ref. [16]. Since the energy density $\rho_{\gamma}(r)$ of photons and pairs is highly peaked near the horizon, as shown in Fig. 2, we use the $\delta$-function approximation $\int(1-r_{+}/r)^{-1/2}\rho_{\gamma}(r)dr\approx\rho_{\gamma}(r_{+})\delta\ell$, and $\delta\ell=(1-r_{+}/r)^{-1/2}dr$ is a proper distance near the horizon. The macroscopic length $\delta\ell$ should be much larger than the Compton length $\lambda_{e}$, but smaller than the fireshell size $2GM\approx 2.33\times 10^{6}$cm.
2.3 Thermodynamic and opaque radiation fluid of photons and pairs
The photons and pairs’ energy (number) density $\rho_{\gamma}$ ($n_{\gamma}$) is extremely large inside the trapped fireshell. The mean free path $\xi_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$ and the thermalization time scale $\tau_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}c$ are much smaller than the macroscopic fireshell size $\bar{R}$ and the time scale $\bar{R}/v_{s}$. Therefore, the local thermal equilibrium is almost instantly established via the back-and-forth process $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$ (2.1) of the cross section,
$\sigma_{\gamma}\approx\sigma_{{}_{T}}\frac{3}{8\omega_{\gamma}}\left[\ln\left(2\omega_{\gamma}\right)-1\right],\quad\omega_{\gamma}=\frac{T_{\gamma}}{m_{e}}\gg 1,$ (2.7)
and time and length scales are
$\tau_{\gamma}=\xi_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}\approx\left[\sigma_{{}_{T}}n_{\gamma}\frac{3}{8\omega_{\gamma}}\left[\ln\left(2\omega_{\gamma}\right)-1\right]\right]^{-1},$ (2.8)
where the Thomson cross section $\sigma_{{}_{T}}=(8\pi/3)\alpha^{2}/m^{2}_{e}$ and the fine structure constant $\alpha=1/137$.
Figure 3: The photon and pair number density $n_{\gamma}$ (left) and mean-free path $\xi_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$ (right) are plotted as functions of radius $r$ in the local observer frame. The Compton number density $m_{e}^{3}=1.74\times 10^{35}/{\rm cm}^{3}$ and length $\lambda_{e}=m_{e}^{-1}=3.86\times 10^{-11}{\rm cm}$.
The local temperature $T_{\gamma}(r)$ represents the mean kinematic energy of photons and electron-positron pairs, which possess zero chemical potential. As a result, the trapped fireshell is an electrically neutral, opaque and thermalised fluid of photons and electron-positron pairs, which we also call a radiation fluid for short. We describe it by a thermodynamic state of thermal temperature $T_{\gamma}(r)$, energy and number densities
$\rho_{\gamma}(r)=\sigma_{{}_{\rm SB}}d_{f}T_{\gamma}^{4}(r),\quad n_{\gamma}(r)=\frac{30\zeta(3)}{\pi^{4}}d_{f}\sigma T_{\gamma}^{3}=3.7d_{f}\sigma_{{}_{\rm SB}}T_{\gamma}^{3}(r),$ (2.9)
in the local frame at the radius $r$. The Stefan-Boltzmann constant $\sigma_{{}_{\rm SB}}=\pi^{2}/15$ for setting the Boltzmann constant $k_{B}=1$. The temperature $T_{\gamma}(r)$ (2.9), Fig. 2 (right), is the averaged kinetic energy of relativistic particles over species $d_{f}$. In the Standard Model (SM) of fundamental particle physics, we adopt $d_{f}\approx 30$ for the number of massless gauge bosons and light fermion-antifermion pairs  [^5] Since the number density is so high, see Fig. 3, neutrinos and anti-neutrinos can be possibly kept inside the fluid in participation of thermalisation., which can kinematically participate in the thermal equilibration at temperature $T_{\gamma}>10$ MeV. The boundary conditions (2.5) gives the temperature $T_{\gamma}^{+}$ near the horizon,
$T_{\gamma}^{+}\equiv T_{\gamma}(r_{+})=\rho^{+}_{\gamma}/(\sigma_{{}_{\rm SB}}d_{f}),$ (2.10)
whose value depends on the balance between ultra-relativistic particles and gravitational pressures. The horizon temperature $T^{+}_{\gamma}$ characterises the maximum temperature of a trapped photon-pair fireshell at the place where the gravitational force is maximum. For these discussions, the equipartition of particle species’ kinetic energies is necessary; an exact thermal equilibrium is not required.
As shown in Figs. 2 and 3, the trapped fireshell’s energy, number densities, and temperature are maximal near the horizon and decrease rapidly in radius. It forms a spherical incandescent halo of ultra-relativistic photons and pairs near the horizon. The halo width is about $2GM=1.16\times 10^{6}$ cm for the temperature varying from the maximum $\sim 10^{2}$ MeV near the horizon to $\sim 1$ MeV near the outer boundary. The halo ends at the outer boundary, where electrons and positrons annihilate to two low-energy (<1<1 MeV) photons, and the reverse process is kinematically impossible. The total number of photons and pairs is enormous
$N_{\gamma}$ $=$ $\int_{r_{+}}^{2r_{+}}n_{\gamma}(r)\frac{4\pi r^{2}dr}{(1-\frac{r_{+}}{r})^{1/2}}\approx n_{\gamma}^{+}4\pi r_{+}^{2}\delta\ell,$ (2.11)
corresponding to the total energy $E_{\gamma}$ (2.6). Here, the maximal photon-pair number density near the horizon is $n_{\gamma}^{+}\equiv n_{\gamma}(r_{+})$, the macroscopic length $\delta\ell$ is approximately equal to the one in Eq. (2.6) for the total energy. Therefore, we effectively describe the trapped fireshells by the temperature $T^{+}_{\gamma}$ and length $\delta\ell$.
The proper time and energy of physical processes occurring in the fireshell near the horizon receive significant gravitational dilation and redshift with respect to an observer at infinity, see Fig. 1 and its caption. The fireshell differs from the photon sphere (ring) of radius $3r_{+}/2$, which describes photons (light) forced to travel in orbits, due to the strong gravitational field near the horizon.
2.4 Temperature radial gradient and anisotropic radiation flux
Photons and pairs in the fireshell are not freely streaming outward, because the mean-free path (2.8) is small, shown in Fig. 3. On the microscopic scale, a local spacetime around each local point $(r,t)$ inside the fireshell, random particle collisions in all directions, the thermal equilibrium state and temperature $T_{\gamma}(r)$ are spatially isotropic without any preferential direction. However, on the macroscopic scale, the temperature $T_{\gamma}(r)$ has a large gradient in the radial direction due to the strong gravitational field near the horizon. Such a temperature $T_{\gamma}(r)$ gradient results in a radially outward diffusion of photons and pairs inside the fireshell. Following the photon-pair energy density $\rho_{\gamma}(r)$ and its value (2.5) near the horizon, we describe the diffusion by the flux $F_{\gamma}=\rho_{\gamma}c$ and luminosity $L_{\gamma}=4\pi r^{2}\rho_{\gamma}c$, whose values near the horizon are
$F_{\gamma}^{+}\equiv\rho_{\gamma}(r_{+})c=\rho_{n}c,~~~L_{\gamma}^{+}\equiv L_{\gamma}(r_{+})=\eta L_{\rm pl},$ (2.12)
where the “Planck luminosity” $L_{\rm pl}\equiv 4\pi r_{+}^{2}\rho_{{}_{M}}c=(3/2)m^{2}_{\rm pl}=3.73\times 10^{59}{\rm ergs/sec}$.
As speculated in Refs. [17, 18] and will be discussed below, the diffusion flux dissipates the fireshell energy $E_{\gamma}$ (2.6) to the blackbody radiation and the emissions of UHE particles. Therefore, the fireshell energy density and pressure $P_{\gamma}=\rho_{\gamma}/3$ decrease. The fireshell is a metastable configuration that gradually shrinks due to gravity. Since the fireshell opacity is large, the energy dissipation rate $\dot{E}_{\gamma}/E_{\gamma}$ is small, and processes are nearly adiabatic. Thus, we will approximately consider the fireshell as a steady-state configuration (source) with a large thermal energy $E_{\gamma}$ (2.6) budget and an enormous number of relativistic particles $N_{\gamma}$ (2.11).
In this fireshell example for highly energetic arenas, we use the TOV equation (2.2) and local temperature $T_{\gamma}(r)$ to describe a macroscopic-sized fireshell of the opaque thermal fluid of dense and energetic photons and pairs, which is trapped around a spherical black hole horizon following a massive stellar core collapse. We ought to mention the following points.
  1. (1)
    One can generalise such a thermal fireshell to macroscopic-sized firespots in strong gravity and non-spherically symmetric environments of astrophysical accretion and coalescence in the presence of angular momentum, magnetic fields and other components.
  2. (2)
    The gravity strength decrease results in the energy density $\rho_{\gamma}$ and temperature $T_{\gamma}$ decreases, the latter is also affected by the number $d_{f}$ (2.9) of particle species participating in the thermal equilibrium. When the temperature (mean kinetic energy) drops below the hadronic binding energy, hadronisation occurs, and the photon-lepton(pair) and gluon-quark(pair) thermal configuration becomes the photon-lepton(pair) one with electrons, protons and neutrons.
As will be discussed below, the dynamics producing UHE particles are on microscopic scales $\sim 10^{-2}$ cm, independently of firespots’ macroscopic shapes and global symmetries. The essentially relevant quantities for the dynamics are maximum temperature $T_{\gamma}^{+}>10$ MeV, the macroscopic size $\delta\ell$ (2.6) and baryon matter contamination $B$ of the firespots.
2.5 Baryon matter contamination, bulk electrons and protons
The trapped fireshell of photons and pairs around the black hole horizon has an ultra-high temperature $T_{\gamma}$, see Fig. 3, which, however, drops quickly below $100$ MeV as the radius increases from the horizon. The quark-gluon plasma undergoes a phase transition of hadronisation and becomes a baryon matter, some of which contaminates the radiation fluid of photons and pairs inside the fireshell. We are interested in the temperature range $T_{\gamma}\sim(10\sim 50)$ MeV, and the baryon matter inside the fireshell is fully ionised, comprising electrons ($e$) and protons ($p$), which are electrically neutral in total.
These electrons ($e$) and protons ($p$) have to be distinguished from electrons $e^{-}$ and positrons $e^{+}$ associated with two photons $2\gamma$ (2.1). They have non-zero chemical potentials and do not participate in the back-and-forth process $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$ (2.1) for the thermal equilibrium of photons and pairs. Therefore, their mean kinetic energies do not share with the mean kinetic energies (temperature $T_{\gamma}$) of photons and pairs. These electrons ($e$) and protons ($p$) effects on the thermal equilibrium, which were studied by using the Boltzmann equation in Ref. [15], are small if their average energy and number densities are much smaller than those of photons and pairs.
Thus, we model such an ionised baryon matter contamination as dilute electron ($e$) gas and proton ($p$) gas immersed inside the thermal and neutral radiation fluid of photons and pairs at the temperature $T_{\gamma}$. In this article, we focus on these electrons ($e$) and protons ($p$), their electromagnetic interaction with a dense radiation fluid of photons and pairs. We ignore neutrons and weak interactions associated with the inverse $\beta$ decay and anti-neutrino production.
Colliding with photons and pairs via multi-particle collisions and field interactions, these electrons ($e$) and protons ($p$) are in microscopic random motions in all directions. They form a relativistic ($e$) gas and a non-relativistic ($p$) gas, and do not form an electron-proton oscillating plasma. We describe them by bulk electrons and protons inside the fireshell, whose numbers are,
$\bar{N}_{e}$ $=$ $\bar{N}_{p}=BN_{\gamma}.$ (2.13)
Here, the baryon-loading $B\ll 1$ quantifies the total number of electrons (protons) inside an opaque fireshell of photons and pairs of the number $N_{\gamma}$. The $B$ should relate to the baryon loading of fireballs, whose outward expansion accounts for GRBs.
We adopt the notations $\tilde{\beta}_{e}$ and $\tilde{\gamma}_{e}$ to represent the velocity and Lorentz factor of bulk electrons  [^6] Bulk electrons (protons) are relativistic (non-relativistic), which indicates individual bulk electron (proton) of Lorentz factor $\tilde{\gamma}_{e}>1$ ($\tilde{\gamma}_{p}\gtrsim 1$) without being averaged over their distribution function in energy-momentum phase space. It is analogous to an opaque radiation fluid of relativistic photons and pairs, whose macroscopic diffusion (drifting) velocities are small, corresponding to an averaged Lorentz factor close to unity., which are in random motions at microscopic scales, therefore,
$\langle\tilde{\beta}_{e}\rangle=0,~~~\langle\tilde{\gamma}_{e}\rangle=1,$ (2.14)
where the averages $\langle\tilde{\beta}_{e}\rangle$ and $\langle\tilde{\gamma}_{e}\rangle$ are over the energy-momentum distribution of bulk electrons. The average values (2.14) mean that bulk electrons have no macroscopic motions in any preferential direction. However, due to an external radiation force or other dynamics at macroscopic scales, the average values can be non-zero: $\langle\tilde{\beta}_{e}\rangle\not=0$ and $\langle\tilde{\gamma}_{e}\rangle>1$. It implies that some of the bulk electrons tend to macroscopic motions in a preferential direction and with a non-vanishing velocity
$\beta_{e}=\langle\tilde{\beta}_{e}\rangle\not=0,~~~\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle>1.$ (2.15)
We call them (2.15) drifting electrons with a non-vanishing velocity $\beta_{e}\not=0$ and Lorentz factor $\gamma_{e}>1$ at macroscopic scales. The same discussions apply to bulk and drifting protons. We now turn to discuss the dynamics of radiation force accelerating electrons from the bulk electrons (2.14) to the drifting electrons (2.15).
3 Compton-rocket effect on accelerating electrons
3.1 Radiation force acting on bulk electrons
Reference [21] studies the Compton-rocket effect: via Thomson scattering, the radially outward photon flux (intensity) $F_{\gamma}(r)=\rho_{\gamma}(r)c$ acts as a radiation force on an optically thin gas of hot electrons or cold protons, and accelerates them to high energies. The photon-scattering excess force on an isotropic gas results from the globally anisotropic loss of internal radiation energy. In the one-dimensional model with respect to a local observer’s proper time $\tau$ in a flat Lorentz frame at the radius $r$, the rates of energy and momentum transfer from the radiation flux $F_{\gamma}(r)$ to the electron gas are (per electron),
$\frac{d\epsilon_{e}}{d\tau}$ $=$ $-\sigma_{{}_{T}}F_{\gamma}(r)[(4/3)(\beta_{e}\gamma_{e})^{2}],$ (3.1)
$\frac{dp_{e}}{d\tau}$ $=$ $\sigma_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{e}\gamma_{e})^{2}+1],$ (3.2)
and some of the bulk electrons [^7] The phrase “some of the bulk electrons” will be quantified by the fraction $N_{e}/\bar{N}_{e}$ (3.4), which equivalently corresponds to the probability of a bulk electron accelerated to the macroscopic motion along the radial direction. are accelerated to drifting electrons with macroscopic velocity $\beta_{e}=v_{e}/c$, Lorentz factor $\gamma_{e}=(1-\beta^{2}_{e})^{1/2}$, and momentum $p_{e}=m_{e}\beta_{e}\gamma_{e}$, $(\beta_{e}\gamma_{e})^{2}=\gamma_{e}^{2}-1$, corresponding to (2.15). The initial conditions at $\tau=0$ for these dynamical equations (3.1) and (3.2) are $\beta_{e}=\langle\tilde{\beta}_{e}\rangle=0$ and $\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle=1$ (2.14), when electrons are bulk electrons. Note that the radiation flux $F_{\gamma}(r)$ due to multi-photon scattering off an electron should be considered as a collective force at macroscopic scales, rather than a single photon scattering off an electron at microscopic scales.
To simplify the notations consistently with the one-dimensional model, we do not specify the averages $\langle\cdot\cdot\cdot\rangle$ over the electron distribution in the energy-momentum phase space. Equation (3.2) gives the Eddington luminosity argument for the incident radiation pressure on non-relativistic electrons ($\beta_{e}\approx 0$) in the Thomson scattering limit [21]. For the proton case, these formulae and discussions remain the same with the substitutions of the subscript $e\rightarrow p$ and approximate cross section $\sigma_{{}_{T}}\rightarrow\sigma^{p}_{{}_{T}}\approx(m_{e}/m_{p})^{2}\sigma_{{}_{T}}\ll\sigma_{{}_{T}}$ of multi-photons and proton collisions.
Radiation fields non-linearly interact with relativistic electrons via the Compton rocket effect [21, 22, 23] by the anisotropy of multi-photons and electron scattering. This dynamics has been studied in an optically thin plasma of electrons and protons by particle-in-cell numerical simulations [24, 25, 26, 27]. It is worthwhile to emphasise that the Compton-rocket effect (3.1,3.2), where the photon flux $F_{\gamma}(r)$ is essentially attributed to the nonlinear QED processes of multi-photon scatterings and interactions (see, e.g., [28, 29, 30]). These non-linear interactions are particularly important when the radiation field is energetic and the charged particle density is large. It is what we will study in the following sections.
3.2 Drifting electrons in an opaque radiation fluid at high temperatures
In the one-dimensional and spherical symmetric fireshell, the bulk electrons averaged momentum $\langle\tilde{p}_{e}\rangle=m_{e}\langle\tilde{\beta}_{e}\tilde{\gamma}_{e}\rangle$ vanishes with the random momentum distribution $\propto e^{-\tilde{p}_{e}^{2}/T_{\gamma}}$ with an equal probability of inward and outward moving. Thus, bulk electrons do not have a net drift motion in the direction $\hat{r}>0$ or opposite one $\hat{r}<0$, and their averaged velocity vanishes $\langle\tilde{\beta}_{e}\rangle=0$ and $\langle\tilde{\gamma}_{e}\rangle=1$ (2.14), and the same discussions for bulk protons $\bar{N}_{p}$.
However, the situation changes when the strong radiation force (3.2) is present, acting on bulk electrons, since the density of photons and pairs has a large variation gradient in the radial direction, as shown in Figs. 2 and 3. We attempt to investigate the Compton-rocket effect on bulk electrons in a highly opaque fluid of photons and pairs at high temperatures $T_{\gamma}\gg m_{e}$, densities $\rho_{\gamma}\gg m_{e}^{4}$ and $n_{\gamma}\gg m_{e}^{3}$. Namely, how the radially outward flux (intensity) $F_{\gamma}(r)=\rho_{\gamma}(r)c\gg m_{e}^{4}$ acts as a radiation force on bulk electrons and protons (2.14), accelerating some of them outward along the radial direction to ultra high energies, drifting out of the fireshell.
This system is more complex than an optically thin electron-proton plasma irradiated at energies $\epsilon_{\gamma}\sim m_{e}$ photons, numerically studied, see e.g. Refs. [27, 31]. Nevertheless, to gain an insight into the dynamics of the Compton-rocket effect on bulk electrons and protons, we use a simplified one-dimensional model in the radial direction only. The photons and pairs flux $F_{\gamma}=\rho_{\gamma}(r)c$ acts on bulk electrons in a tiny tube with the cross section $\sigma_{\gamma}$.
3.3 Drifting electrons versus bulk electrons
Although the mean-field path of bulk electrons is small, as shown in Fig. 3, the outward radiation force $F_{\gamma}=\rho_{\gamma}(r)c$ is strong, so some bulk electrons have chances within their mean-free path to get accelerated and drift radially outwards. Therefore, some of bulk electrons drift in the direction $\hat{r}>0$ and their averaged momentum $p_{e}=\langle\tilde{p}_{e}\rangle_{a}$ and energy $\epsilon_{e}=(p_{e}^{2}+m_{e}^{2})^{1/2}$ do not vanish. They are drifting electrons with the macroscopic velocity $\beta_{e}$ and Lorentz factor $\gamma_{e}$ (2.15), and we define the number of these outward drifting electrons as
$N_{e}(\gamma_{e})<\bar{N}_{e},~~~\beta_{e}=\langle\tilde{\beta}_{e}\rangle_{a}>0,~~~\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle_{a}>1,$ (3.3)
where the average $\langle\cdot\cdot\cdot\rangle_{a}$ is over the energy-momentum distribution of bulk and drifting electrons. The drifting electrons $N_{e}(\gamma_{e})$ should be a small fraction of the bulk electrons $\bar{N}_{e}$, because most of the radiation force accelerated electrons do not become drifting electrons, but run back to bulk electrons due to collisions with photons and pairs. We define the fraction (radio) of drifting and bulk electrons as
$\frac{N_{e}}{\bar{N}_{e}}\equiv\frac{N_{e}(\gamma_{e})}{\bar{N}_{e}},$ (3.4)
which actually represents the probability that electrons are accelerated out of bulk electrons to become drifting electrons by the Compton-rocket effect in an opaque photon-pair fluid. This fraction (3.4) is a crucial quantity we adopt to describe the significance of the Compton-rocket effect on accelerating electrons to a Lorentz factor $\gamma_{e}$ in an opaque photon-pair fluid. The same discussions apply to drifting $N_{p}$ and bulk $\bar{N}_{p}$ protons in the opaque fluid of photons and pairs.
We further examine the time scale $\tau_{a}^{e}$ of the electron acceleration process (Compton-rocket effect) and compare it with the thermalisation time scale $\tau_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}c$ (2.8) of photon and pair fluid. Using Eq. (3.2) we define
$\tau_{a}^{e}\equiv\frac{p_{e}}{dp_{e}/d\tau}=\frac{m_{e}\beta_{e}\gamma_{e}}{\sigma_{{}_{T}}F_{\gamma}[\frac{2}{3}(\beta_{e}\gamma_{e})^{2}+1]},$ (3.5)
and obtain the ratio
$\frac{\tau_{a}^{e}}{\tau_{\gamma}}=\frac{3.7\times 9}{16\omega^{2}_{\gamma}}\frac{\ln\left(2\omega_{\gamma}\right)-1}{(\beta_{e}\gamma_{e})\left[1+\frac{3}{2(\beta_{e}\gamma_{e})^{2}}\right]}.$ (3.6)
Here, the gravitational time dilation factors $d\tau=(1-1/r)^{1/2}dt$ are simplified. Recall that the thermalisation time scale $\tau_{\gamma}$ (2.8) represents also the time scale of the radiation flux $F_{\gamma}$ of multi-photons acting on a changed particle.
The time and length scales $\tau_{a}^{e}$ of the Compton-rocket effect are smaller than the thermalisation timescale $\tau_{\gamma}$ of photons and pairs fluid, i.e., $\tau_{a}^{e}<\tau_{\gamma}$, for $\omega_{\gamma}=T_{\gamma}/m_{e}>1$ and $\gamma_{e}>1$. The drifting electron density $n_{e}(\gamma_{e})$ is much smaller than the density $n_{\gamma}$ of photons and pairs. These properties imply that drifting electrons accelerated by the Compton-rocket effect are small and rapid perturbations upon the thermal bath of photons and pairs at high energies and densities. If the timescale inequality was $\tau_{a}^{e}>\tau_{\gamma}$ instead of $\tau_{a}^{e}<\tau_{\gamma}$, the perturbations of drifting electrons would have been washed out by photon-pair thermal fluctuations. The drifting electron perturbation’s energy and momentum are $\delta\epsilon_{e}>m_{e}$ and $\delta p_{e}>m_{e}$ with Lorentz factor $\gamma_{e}\gtrsim 1$.
Moreover, due to the large opacity and frequent collisions with photons and pairs, drifting electrons ($\gamma_{e}\gtrsim 1$) quickly dissipate their energy and run back to bulk electrons ($\gamma_{e}=1$), which are in an energy equipartition with photons and pairs. It implies that the perturbations of drifting electrons ($\gamma_{e}>1$) should be rapidly washed out. Therefore, it looks unlikely that in the opaque fluid of photons and pairs, such small perturbations could develop, namely, some accelerated electrons by the Compton-rocket effect could drift to a large Lorentz factor $\gamma_{e}\gg 1$, unless some instabilities of avalanche phenomena occur.
4 Electrons runaway in an opaque radiation fluid
4.1 Instability and electron runaway due to Klein–Nishina corrections
As discussed, in the opaque fluid of photons and pairs, the Compton-rocket effect on accelerating electrons is an approximately “local” and “instantaneous” perturbation that triggers electrons to increase energy and drift in the outward radial direction. On their way, accelerated electrons collide with photons and pairs with the probability described by the Klein–Nishina (KN) cross section (4.1), which, however, decreases as the accelerated electrons’ energy increases, and the straightforward scattering with a narrower deflection angle $\theta_{e}\sim 0$. This property of KN corrections implies the instability of an unstable avalanche runaway process of drifting electrons. Namely, accelerated electrons gain more energy and have fewer collisions. Thus, they have greater chances to drift a longer distance and gain even more energy from the Compton-rocket effect. As a result, there is a probability that a small fraction of drifting electrons run away and achieve ultra-high energies, penetrating through the opaque fluid of photons and pairs.
In the opaque fluid of photons and pairs at the temperature $T_{\gamma}>m_{e}$, drifting electrons move outward in the radial direction $\theta_{e}\sim 0$, with the velocity $\beta_{e}\sim 1$ and Lorentz factor $\gamma_{e}>1$. The KN cross-section is approximately given by
$\sigma^{e}_{{}_{\rm KN}}\approx\sigma_{{}_{T}}\frac{3}{8\omega_{e}}\left[\ln\left(2\omega_{e}\right)+\frac{1}{2}\right],\quad\omega_{e}\approx\frac{2T_{\gamma}}{m_{e}}\gamma_{e},$ (4.1)
where the factor $2$ in $\omega_{e}$ (4.1) is due to the blue-shift effect [32]. The cross-section decreases as the drifting electrons’ Lorentz factor $\gamma_{e}$ increases and vanishes when $\gamma_{e}\gg 1$. Therefore, high $\gamma_{e}$ drifting electrons have a small probability of being scattered back to bulk electrons with $\gamma_{e}=1$. Instead, they are further accelerated to higher $\gamma_{e}$ by the Compton-rocket effect, and behave as runaway drifting electrons.
Such a runaway process has a time scale $\tau^{e}_{{}_{\rm KN}}=(\sigma^{e}_{{}_{\rm KN}}n_{\gamma}c)^{-1}$ of the instability, which is much larger than the thermalization time scale $\tau_{\gamma}$ (2.8) of photons and pairs,
$\frac{\tau^{e}_{{}_{\rm KN}}}{\tau_{\gamma}}\approx 2\gamma_{e}\frac{\ln\left(2T_{\gamma}/m_{e}\right)}{\ln\left(4\gamma_{e}T_{\gamma}/m_{e}\right)}\gg 1,~~~\gamma_{e}\gg 1,$ (4.2)
and the gravitational time dilation factors $d\tau=(1-1/r)^{1/2}dt$ are simplified. The time scales (3.5), (3.6) and (4.2) give $\tau^{e}_{a}<\tau_{\gamma}\ll\tau^{e}_{{}_{\rm KN}}$, implying that rather than being smeared out, the instability could develop a runaway solution. It makes drifting electrons start from microscopic perturbations (3.6) and run to macroscopic motions (4.2).
4.2 Probability of drifting electrons runaway
Even if such an instability occurs, in view of drifting electrons colliding with photons and pairs, what is the probability that runaway drifting electrons survive when they travel over a long macroscopic distance and time scale, reaching a large Lorentz factor $\gamma_{e}$? We describe this probability by the fraction $N_{e}/\bar{N}_{e}$ (3.4) of drifting and bulk electron numbers, and study it for the following three cases:
  1. (i)
    $N_{e}/\bar{N}_{e}=1$ (maximal probability) corresponds to all of the bulk electrons being accelerated and becoming drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. This collisionless case can not be realistic, since the drifting electrons dissipate their kinetic energies, and return to bulk electrons via collisions with photons and pairs in the fireshell of a large opacity. Nevertheless, we study this case to gain an insight into the nature of the Compton-rocket effect on accelerating electrons inside an opaque fluid of energetic and dense photons and pairs.
  2. (ii)
    $N_{e}/\bar{N}_{e}\approx 0$ (near zero probability) corresponds to almost none of the bulk electrons being accelerated and becoming drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. In other words, almost all drifting electrons return to bulk electrons due to collisions with photons and pairs. It indicates that the instability does not develop into an avalanche runaway process, and a negligible amount of bulk electrons are accelerated and become drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. It is the collisional case of Thomson scattering with an energy-independent cross section at low energies.
  3. (iii)
    $N_{e}/\bar{N}_{e}\not=0$ is a non-trivial value (finite probabilities), corresponding to a significant amount of the bulk electrons being accelerated and becoming drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. It indicates that the instability has developed into an avalanche runaway process, and a significant amount of bulk electrons are accelerated and become drifting electrons at Lorentz factor $\gamma_{e}\gg 1$. It is the collisional case of Klein–Nishina scattering with an energy-dependent cross section at high energies.
To quantitatively study this fraction $N_{e}/\bar{N}_{e}$ in these cases, we need two master equations for drifting electrons: (1) the acceleration equation (4.3) attributed to the Compton-rocket effect, (2) the decay rate equation (4.8) or (4.13) due to drifting electrons’ energy-momentum dissipation in collisions with photons and pairs inside the fireshell.
We rewrite the Compton-rocket effect on the electron acceleration (3.2) in a small distance $dr=\beta_{e}d\tau=\beta_{e}(1-r_{+}/r)^{1/2}dt$ as
$m^{\rm eff}_{e}\gamma_{e}\frac{d\gamma_{e}}{dr}$ $=$ $\frac{1}{3}\sigma_{{}_{T}}(1-r_{+}/r)^{-1/2}F_{\gamma}(r)(2\gamma_{e}^{2}+1).$ (4.3)
In the one-dimensional model, this equation represents drifting electrons $N_{e}(\gamma_{e})$ of $\gamma_{e}>1$ accelerated out of bulk electrons $\bar{N}_{e}$ of $\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle=1$. The boundary condition is $\gamma_{e}=1$ at $r=r_{+}$ from the initial condition for (3.2). The effective mass $m^{\rm eff}_{e}\approx m_{p}+m_{e}$ will be explained in due course.
4.3 Electrons accelerated by an opaque radiation fluid at high temperature
To have an idea about the spacetime length scales of electron acceleration in the fireshell of energetic and dense photons and pairs, we first consider the collisionless case where all bulk electrons $\bar{N}_{e}$ (2.13) are accelerated to drifting electrons $N_{e}(\gamma_{e})$ in the radial direction without any collision with photons and pairs. Then integrating the Compton-rocket effect (4.3) yields
$2\gamma_{e}^{2}+1=3\exp\Big[\frac{4\sigma_{{}_{T}}}{3m^{\rm eff}_{e}}\int_{r_{+}}^{r}\frac{\rho_{\gamma}(r^{\prime})}{(1-r_{+}/r^{\prime})^{1/2}}dr^{\prime}\Big],$ (4.4)
where bulk electrons $\gamma_{e}=1$ near the horizon $r_{+}$ are accelerated to drifting electrons $\gamma_{e}>1$ at the radius $r$. The photon and pair flux $F_{\gamma}(r)=\rho_{\gamma}c$ is highly peaked and has a great gradient near the horizon $r_{+}$, as shown in Fig. 2. Therefore, the photon and pair flux $F^{+}_{\gamma}=\rho_{\gamma}(r_{+})c$ (2.12) and temperature $T^{+}_{\gamma}=T_{\gamma}(r_{+})$ (2.10) near the horizon predominantly contribute to the integration (4.4), i.e., to the acceleration of drifting electrons out of bulk electrons. Similarly to Eq. (2.6) for the total energy, we use the $\delta$-function approximation $\int(1-r_{+}/r)^{-1/2}\rho_{\gamma}(r)dr\approx\rho_{\gamma}(r_{+})\delta\ell=\sigma_{{}_{\rm SB}}d_{f}T^{4}_{\gamma}(r_{+})\delta\ell$, where $\delta\ell=(1-r_{+}/r)^{-1/2}dr$ is a small proper distance near the horizon. Solution (4.4) approximately becomes
$2\gamma_{e}^{2}+1\approx 3\exp\Big[\frac{4\sigma_{{}_{T}}\sigma_{{}_{\rm SB}}d_{f}}{3m^{\rm eff}_{e}m_{e}}T^{4}_{\gamma}(r_{+})\delta\ell\Big],$ (4.5)
namely, the radiation force $F_{\gamma}(r_{+})=\rho_{\gamma}(r_{+})c$ near the horizon, where the energy density $\rho_{\gamma}(r_{+})$ and temperature $T^{+}_{\gamma}$ are maximal, accelerates on an electron over a distance $\delta\ell$ to reach its Lorentz factor $\gamma_{e}$.
In the viewpoint of the energy density $\rho_{\gamma}\sim T_{\gamma}^{4}\sim E^{2}_{\rm eff}/(8\pi)$, the photons and pairs radiation force (pressure) $F_{\gamma}=c\rho_{\gamma}$ for the temperature $T_{\gamma}^{+}\equiv T_{\gamma}(r_{+})\sim{\mathcal{O}}(10^{1-2})$MeV is equivalent to an effective electric force $eE_{\rm eff}$, which is about ${\mathcal{O}}(10^{1-2})$ times larger than the critical electric field force $eE_{c}=m_{e}^{2}$ for electron-positron pair productions. In a classical scenario of a test charged particle accelerated by an electric field, we estimated that such an effective electric force $eE_{\rm eff}$ accelerates an electron to PeV $\sim 10^{9}m_{e}(\gamma_{e}\sim 10^{9})$ energies in a small distance $d\sim{\mathcal{O}}(10^{9})\lambda_{e}\sim{\mathcal{O}}(10^{-2})$ cm and time $\tau=d/c\sim{\mathcal{O}}(10^{-12})$ sec [17]. Therefore, the macroscopic length $\delta\ell$ determined from the total thermal energy $E_{\gamma}$ (2.11) should be large enough for an electron reaching $\gamma_{e}\sim 10^{9}$ in Eq. (4.5). It indicates that the macroscopic sizes and geometries of dense and energetic photon-pair fireshells or spots are irrelevant for the Compton-rocket effect, because the local inhomogeneous density of photons and pairs accelerates electrons on microscopic time and length scales $d$, which are much smaller than the macroscopic time and size variation of fireshells or firespots.
Physically, the collisionless solution (4.5) means that (i) all bulk electrons $\bar{N}_{e}$ of $\gamma_{e}=1$ accelerated to drifting electrons $N_{e}(\gamma_{e})$ of an enormous Lorentz factor $\gamma_{e}\gg 1$; (ii) the majority of drifting electrons get accelerated and reach their ultra-high energies in the neighbourhood near the horizon, or where photon-pair energy density and temperature are maximal. This collisionless solution (4.5) of $N_{e}(\gamma_{e})=\bar{N}_{e}$ is certainly not realistic, since the radiation fluid of photons and pairs is extremely opaque, and drifting electrons colliding with photons and pairs is unavailable. However, the study of the collisionless case (4.5) let us learn (i) the large-temperature $T^{+}_{\gamma}$ dominates the radiation-force contribution to electron acceleration; (ii) the spacetime length scales $d$ of electron acceleration in an opaque radiation fluid of photons and pairs. This helps make appropriate approximations in studying collisional cases, in which the drifting electron cannot reach $\gamma_{e}\gg 1$ (4.5) with the maximal probability (fraction) $N_{e}/\bar{N}_{e}=1$.
For the collisional case, to show the runaway dynamics due to the Klein-Nishina correction in high energies, as previously speculated in section (4.1), we attempt to calculate the fraction $N_{e}/\bar{N}_{e}$ numerically in two separate cases: (1) the stable case, where drifting electrons undergo Thomson scattering in low energies; (2) the unstable case, where drifting electrons undergo Klein-Nishina scattering in high energies. We do not adopt the full scattering cross-section that varies with energy. The reasons are that (i) there is no analytical solution $N_{e}/\bar{N}_{e}$ for the entire energy range; (ii) we have numerical difficulties to tackle the runway dynamics when it occurs, the fraction $N_{e}/\bar{N}_{e}$ varies by many orders of magnitude, as will be shown in Figs. 4 and 5; (iii) it is helpful to distinctly compare and contrast Thomson and Klein-Nishina scattering cases to understand how the runaway dynamics physically develops.
4.4 Stable case: Thomson scattering at low energies
Second, we consider the collisional case of the constant Thomason scattering cross section $\sigma_{{}_{T}}$ at low energies (temperatures). We calculate the number of drifting electrons accelerated in one proper mean-free path $\xi_{e}^{\rm T}=(n_{\gamma}\sigma_{{}_{T}})^{-1}$ and time $d\tau=\xi_{e}^{\rm T}/\beta_{e}$. The acceleration equation (3.2) or (4.3) for the Lorentz factor variation $\Delta\gamma^{\rm T}_{e}$ in one mean-free path becomes
$m^{\rm eff}_{e}\gamma_{e}\Delta\gamma^{\rm T}_{e}$ $=$ $\frac{1}{3}\sigma_{{}_{T}}F_{\gamma}(r)(2\gamma_{e}^{2}+1)\xi_{e}^{\rm T}.$ (4.6)
Using the relation $F_{\gamma}/n_{\gamma}=T_{\gamma}/3.7$, the variation $\Delta\gamma^{\rm T}_{e}$ in one mean-free path is given by
$\Delta\gamma^{\rm T}_{e}$ $=$ $\frac{1}{3.7}\left(\frac{T_{\gamma}}{m^{\rm eff}_{e}}\right)\frac{2\gamma_{e}^{2}+1}{3\gamma_{e}},$ (4.7)
which is proportional to $(T_{\gamma}/m^{\rm eff}_{e})\gamma_{e}$ for $\gamma_{e}\gg 1$.
Due to the Thomson scatterings with photons and pairs, the drifting electron number $N_{e}(\gamma_{e})$ decreases following the one-dimensional rate equation,
$\frac{dN_{e}}{dr}$ $=$ $-\frac{1}{\xi_{e}^{\rm T}}N_{e}=-n_{\gamma}\sigma_{{}_{T}}N_{e},$ (4.8)
similarly to the equation for unstable particles of the number $N_{e}$ decaying while travelling along the radial trajectory. The rate equation (4.8) recasts as
$\frac{dN_{e}(\gamma_{e})}{d\gamma_{e}}$ $=$ $-\frac{1}{\Delta\gamma^{\rm T}_{e}}N_{e}(\gamma_{e}),$ (4.9)
by using Eqs. (4.3) and (4.6).
Assuming that the majority of drifting electrons get accelerated in the neighbourhood near the horizon, as discussed in Eq. (4.5), we approximate the temperature $T_{\gamma}\approx T^{+}_{\gamma}$ by the horizon temperature $T^{+}_{\gamma}$. Integrating the rate equation (4.9) shows the fraction (probability) of drifting electrons $N_{e}(\gamma_{e})$ out of bulk electrons $\bar{N}_{e}$ ($\gamma_{e}=1$) is given by
$\frac{N_{e}(\gamma_{e})}{\bar{N}_{e}}$ $=$ $\exp\left(-\int_{1}^{\gamma_{e}}\frac{1}{\Delta\gamma^{\rm T}_{e}}{d\gamma_{e}}\right)\approx\left(\frac{3}{2\gamma_{e}^{2}+1}\right)^{2.78(m^{\rm eff}_{e}/T^{+}_{\gamma})}$. (4.10)
As shown on the left of Fig. 4, the fraction $N_{e}(\gamma_{e})/\bar{N}_{e}$ decreases rapidly as the accelerated electron Lorentz factor $\gamma_{e}$ increases on the way they travel outward, dissipating their energy and momentum into the fluid via collisions with photons and pairs due to the large opacity.
Thus, the most accelerated drifting electrons $N_{e}(\gamma_{e})$ at high energies run back to the bulk electrons $\bar{N}_{e}$ at low energies, which are in the energy equipartition with photons and pairs. In other words, the probability of drifting electrons reaching a large Lorentz factor $\gamma_{e}\gg 1$ goes to zero at low temperature $T^{+}_{\gamma}\ll m^{\rm eff}_{e}$. Very few electrons are in the high-energy superthermal spectrum tail; thus, they are negligible and irrelevant for observations of UHE particles. It is the case for the Sun, whose core temperature of photons is much smaller than the electron mass, i.e., $T_{\odot}\approx 1.5\times 10^{4}{\rm eV}\ll m_{e}$ and the fraction (4.10) is exponentially suppressed.
Figure 4: We calculate the fraction (probability) $N_{e}/\bar{N}_{e}$ of accelerated electrons drifting out of the bulk electrons for the Thomson (4.10) and the Klein-Nishina (4.15) scattering cases. It shows that the probability of accelerated electrons reaching high energies is non-trivial, implying the runaway behaviour, in the Klein-Nishina scattering case, but is completely negligible in the Thomson scattering case. These figures also represent the proton case with the substitution of the subscript $e\rightarrow p$.
4.5 Unstable case: Klein-Nishina scattering at high energies
Third, we investigate the collisional case of Klein-Nishina scattering at high energies (temperatures), and show a non-trivial probability of drifting electrons running away, reaching $\gamma_{e}\gg 1$, due to the energy-dependent nature of the KN cross section $\sigma^{e}_{{}_{\rm KN}}(\gamma_{e})$ (4.1), which decreases with increasing energies.
Analogously to $\Delta\gamma^{\rm T}_{e}$ (4.6), define $\Delta\gamma^{\rm KN}_{e}$ to be the Lorentz factor variation of drifting electrons accelerated in one proper mean-free path $\xi_{e}^{\rm KN}=(n_{\gamma}\sigma^{e}_{{}_{\rm KN}})^{-1}$. The acceleration equation (3.2) or (4.3) for the Lorentz factor variation $\Delta\gamma^{\rm KN}_{e}$ in one mean-free path $\xi_{e}^{\rm KN}$ becomes
$m^{\rm eff}_{e}\gamma_{e}\Delta\gamma^{\rm KN}_{e}$ $=$ $\frac{1}{3}\sigma_{{}_{T}}F_{\gamma}(r)(2\gamma_{e}^{2}+1)\xi_{e}^{\rm KN}.$ (4.11)
Using the relation $F_{\gamma}/n_{\gamma}=T_{\gamma}/3.7$ and the KN cross section (4.1) that depends on drifting electrons’ energy, i.e., Lorentz factor $\gamma_{e}$, we obtain the variation $\Delta\gamma^{\rm KN}_{e}$ in one mean-free path given by
$\Delta\gamma^{\rm KN}_{e}(T_{\gamma},\gamma_{e})$ $=$ $\frac{1}{3.7}\frac{16}{9}\left(\frac{T^{2}_{\gamma}}{m_{e}m^{\rm eff}_{e}}\right)\frac{2\gamma_{e}^{2}+1}{\ln(4\gamma_{e}\frac{T_{\gamma}}{m_{e}})+\frac{1}{2}},$ (4.12)
in the contrast with $\Delta\gamma^{\rm T}_{e}$ (4.7) for the Thomason scattering case.
Analogously to Eqs. (4.8) and (4.9) in the Thomson scattering case, due to the Klein-Nishina scatterings with photons and pairs, the drifting electron number $N_{e}(\gamma_{e})$ decreases following the one-dimensional rate equation
$\frac{dN_{e}}{dr}$ $=$ $-\frac{1}{\xi^{\rm KN}_{e}}N_{e}=-n_{\gamma}\sigma^{e}_{{}_{\rm KN}}N_{e}.$ (4.13)
This rate equation is recast as
$\frac{dN_{e}(T_{\gamma},\gamma_{e})}{d\gamma_{e}}$ $=$ $-\frac{1}{\Delta\gamma^{\rm KN}_{e}(T_{\gamma},\gamma_{e})}N_{e}(T_{\gamma},\gamma_{e}),$ (4.14)
by using Eqs. (4.3) and (4.11).
The integration of the rate equation (4.14) leads to the fraction (probability) of drifting electrons $N_{e}(T_{\gamma},\gamma_{e}>1)$ out of bulk electrons $\bar{N}_{e}(T_{\gamma},\gamma_{e}=1)$,
$\frac{N_{e}}{\bar{N_{e}}}\equiv\frac{N_{e}(T_{\gamma},\gamma_{e}>1)}{\bar{N}_{e}(T_{\gamma},\gamma_{e}=1)}$ $\approx$ $\exp-\int_{1}^{\gamma_{e}}\frac{1}{\Delta\gamma^{\rm KN}_{e}(T^{+}_{\gamma},\gamma^{\prime}_{e})}{d\gamma^{\prime}_{e}}.$ (4.15)
Here, we approximate the radial-dependent temperature $T_{\gamma}(r)$ by the characteristic temperature $T^{+}_{\gamma}$ for the same arguments discussed in Eqs. (4.5) and (4.10). Figure 5 gives the numerical integration (4.15) that shows the non-trivial fraction $N_{e}/\bar{N}_{e}$ of drifting electrons at ultra-high energies, which is in distinct contrast with the completely negligible one (4.10) in the Thomson scattering case. This non-trivial number $N_{e}$ of drifting electrons ($\gamma_{e}\gg 1$) means a significant superthermal spectrum in the high-energy tail, which could be relevant for observations of UHE particles.
Figure 5: We show the fraction $N_{e}/\bar{N}_{e}$ (4.15) of accelerated electrons drifting out of bulk electrons at different temperatures $T^{+}_{\gamma}\gg m_{e}$ near the horizon. The $N_{e}/\bar{N}_{e}$ value is exponentially sensitive to the horizon temperature $T_{\gamma}^{+}$. The probability of accelerated electrons reaching high energies is small for temperatures $T^{+}_{\gamma}\lesssim 10m_{e}$, and the Compton-rocket effect and runaway dynamics might not be observationally relevant. These figures also represent the proton case by substituting the subscript $e\rightarrow p$.
For different temperatures $T^{+}_{\gamma}\approx 100m_{e}\approx 50$ MeV and $T^{+}_{\gamma}\approx 20m_{e}\approx 10$ MeV, Figure 5 shows the following spectral features of the probability that electrons get accelerated to a large Lorentz $\gamma_{e}$ factor:
  1. (a)
    for Lorentz factors $\gamma_{e}<10^{3}$, the accelerated electron fraction $N_{e}/\bar{N}_{e}$ decreases as $\gamma_{e}$ increases, which may be approximated as a decay power law
    $\frac{N_{e}}{\bar{N}_{e}}\propto(\gamma_{e})^{-\nu},~~\nu>0,~~(\gamma_{e}<10^{3});$ (4.16)
  2. (b)
    for Lorentz factors $\gamma_{e}>10^{3}$, the fraction approach a constancy
    $\frac{N_{e}}{\bar{N}_{e}}~\propto(\gamma_{e})^{0},~~~(\gamma_{e}>10^{3}).$ (4.17)
The feature (a) is expected in a usual physical situation where the probability of drifting electrons running to high energies becomes smaller, due to their kinetic energy dissipations in collisions with photons and pairs. However, the feature (b) is peculiar, showing that the probability $N_{e}/\bar{N}_{e}$ does not decrease as (a), instead, becomes very flat when $\gamma_{e}>10^{3}$. These are the exact attributes of unstable runaway behaviour. Since the KN cross-section decreases as the drifting electrons’ energies increase, they continue accelerating to higher and higher energies by the Compton-rocket effect without dissipating much energy into photons and pairs in the radiation fluid. Such an avalanche dynamics is crucially important for producing UHE particles of Lorentz factors $\gamma_{e}\gg 1$.
As the temperature $T_{\gamma}$ decreases and approaches $2m_{e}$, the asymptotic value $N_{e}/\bar{N}_{e}$ (4.17) is exponentially suppressed by many orders of magnitude. Some technical difficulties occur in making numerical calculations. Therefore, using $\Delta\gamma^{\rm KN}_{e}$ (4.12) and $(N_{e}/\bar{N}_{e})$ (4.15), we approximately calculate the asymptotic constancy (4.17) for large $\gamma_{e}>10^{3}$
$\left(\frac{N_{e}}{\bar{N}_{e}}\right)_{\gamma_{e}\gg 1}$ $\approx$ $\left(\frac{4T^{+}_{\gamma}}{m_{e}}\right)^{\chi}e^{-\chi},\quad\chi=\frac{m_{p}}{m_{e}}\left(\frac{m_{e}}{T^{+}_{\gamma}}\right)^{2},$ (4.18)
which is exponentially suppressed by small horizon temperatures $T^{+}_{\gamma}/m_{e}$, as can be seen in Figs. 4 and 5. This result (4.18) implies the following features. (i) The significantly nontrivial fraction $N_{e}/\bar{N}_{e}$ of high-energy electrons ($\gamma_{e}\gg 1$) can only be produced in opaque fireshells or spots of the radiation fluid of photons and pairs at temperatures $T_{\gamma}>10m_{e}$, which can be present in highly energetic arenas of astrophysical sources and the early Universe. (ii) In the case of the Sun ($T_{\odot}/m_{e}\ll 1$), the excess of large $\gamma_{e}$ electrons $N_{e}$ in the superthermal tail deviating from the blackbody radiation spectrum is too small to be relevant for observations. (iii) However, the probability of $\gamma_{e}\gg 1$ electrons’ excess may be relevant for the detections in future laser experiments [33] for the physics of strong radiation field intensities $I_{\gamma}>E^{2}_{c}/(8\pi)$, equivalently to high temperature $T^{+}_{\gamma}>10m_{e}$.
To close this section, we mention that for the optically thin system of electron-proton plasma interacting with photons at energies $\epsilon_{\gamma}\sim m_{e}$, Equations of types (4.8,4.9) or (4.13,4.14) were used in the numerical studies [27]. Their results show the superthermal spectrum tail of runaway electrons $N_{e}$ drifting backwards to massive protons due to the electric field induced by the charge separation between electrons and protons.Here, instead, we study electrons and protons that are fully ionised and immersed inside an opaque radiation fluid of photons and pairs. They interact with photons and pairs at very high temperatures $T_{\gamma}\gg m_{e}$, energy density $\rho_{\gamma}\gg m_{e}^{4}$ and number density $n_{\gamma}\gg m_{e}^{3}$.
5 Runaway protons pulled by runaway electrons
5.1 Electric field between drifting electrons and protons
A non-trivial amount of high-energy electrons running away indicates that we cannot ignore protons in the dynamics of the Compton-rocket effect on electrons, because of the electric field $E_{pe}$ developed between separated electrons and protons. This is due to the very different cross sections $\sigma_{{}_{T}}\gg\sigma^{p}_{{}_{T}}\approx(m_{e}/m_{p})^{2}\sigma_{{}_{T}}$ of photons scattering with electrons and protons. In the one-dimensional model, the Compton-rocket effect on protons ($p_{p}=m_{p}\beta_{p}\gamma_{p}$)
$\frac{d\epsilon_{p}}{d\tau}$ $=$ $-\sigma^{p}_{{}_{T}}F_{\gamma}(r)[(4/3)(\beta_{p}\gamma_{p})^{2}],$ (5.1)
$\frac{dp_{p}}{d\tau}$ $=$ $\sigma^{p}_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{p}\gamma_{p})^{2}+1],$ (5.2)
is much smaller than the one (3.2) for electrons.
As a result, charge separation and the electric field $E_{pe}$ between electrons and protons occur. It must back-react on the one-dimensional acceleration equations for both electrons (3.2) and protons (5.2),
$\frac{dp_{e}}{d\tau}$ $=$ $\sigma_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{e}\gamma_{e})^{2}+1]-eE_{pe}>0,$ (5.3)
$\frac{dp_{p}}{d\tau}$ $=$ $\sigma^{p}_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{p}\gamma_{p})^{2}+1]+eE_{pe}\approx eE_{pe}>0.$ (5.4)
It is a highly non-linear dynamical system. We speculate the dynamical condition when photons and pairs’ radiation force $\sigma_{{}_{T}}F_{\gamma}\propto T_{\gamma}^{4}(\beta_{e}\gamma_{e})^{2}$ in Eq. (5.3) is so strong that
  1. (i)
    a significantly non-trivial fraction $N_{e}/\bar{N}_{e}$ of accelerated electrons ($\gamma_{e}\gg 1$) drifting outward, i.e., $dp_{e}/d\tau>0$ in Eq. (5.3) for electrons;
  2. (ii)
    it thus induces a strong and local electric field $E_{pe}=eN_{e}/(4\pi r^{2})$, which is an electric field between a thin spherical capacitor of electron and proton layers in the one-dimensional model;
  3. (iii)
    this electric field binds electron and positron together, pulls massive protons $dp_{p}/d\tau\approx eE_{pe}>0$ in Eq. (5.4), where the radiation force term $\propto\sigma^{p}_{{}_{T}}$ in its right-handed side is negligible. As a result, a significantly non-trivial fraction $N_{p}/\bar{N}_{p}$ of accelerated protons ($\gamma_{p}\gg 1$) drift outwards in the radial direction, following the accelerated electrons’ fraction $N_{e}/\bar{N}_{e}$.
5.2 Strong radiation field accelerating electrons and protons
The bulk and drifting electron numbers are identical to the bulk and drifting proton numbers
$\bar{N}_{e}=\bar{N}_{p},~~~N_{e}=N_{p},~~~N_{e}/\bar{N}_{e}=N_{p}/\bar{N}_{p},$ (5.5)
because of the global charge neutrality of ionised electrons and protons, which are immersed inside an electrically neutral radiation fluid of photons and pairs. The combination of Eqs. (5.3) and (5.4) yields the one-dimensional acceleration equation for electrons and protons together
$\frac{d(p_{e}+p_{p})}{d\tau}$ $\approx$ $\sigma_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{e}\gamma_{e})^{2}+1],$ (5.6)
neglecting the radiation force on protons. Under the condition of a strong enough radiation flux $F_{\gamma}$ and the electric field $E_{pe}$ between electrons and protons, we assume that drifting electrons $N_{e}$ and protons $N_{p}$ are moving together
$\beta_{e}\approx\beta_{p}~~~{\rm and}~~~\gamma_{e}\approx\gamma_{p},$ (5.7)
through the radiation fluid of photons and pairs. The acceleration equation (5.6) is simplified as
$(m_{e}+m_{p})\frac{d(\beta_{e}\gamma_{e})}{d\tau}$ $\approx$ $\sigma_{{}_{T}}F_{\gamma}(r)[(2/3)(\beta_{e}\gamma_{e})^{2}+1].$ (5.8)
This is the same as the electron acceleration equation (3.2) or (4.3) by introducing an effective electron mass
$m^{\rm eff}_{e}=m_{p}+m_{e}\approx m_{p},$ (5.9)
which previously appeared in Eq. (4.3) without explanations. Similarly, the combination of (3.1) and (5.1) yields the energy-gain equation for accelerated electrons and protons
$\frac{d\epsilon_{ep}}{d\tau}$ $\approx$ $-\sigma_{{}_{T}}F_{\gamma}(r)[(4/3)(\beta_{e}\gamma_{e})^{2}],~~~\epsilon_{ep}=\epsilon_{e}+\epsilon_{p},$ (5.10)
where $\epsilon_{e}=m_{e}\gamma_{e}$ and $\epsilon_{p}=m_{p}\gamma_{p}$.
Equations (5.8) and (5.10) qualitatively describe an accelerating system of electrons and protons driven by the radiation force, assuming electrons and protons are tightly associated by their own electric field and accelerated to the approximately same Lorentz factor $\gamma_{e}\approx\gamma_{p}$. In this approximation, all dynamical equations remain the same as those for electrons, except for replacing the electron mass $m_{e}$ with the effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}$ (5.9) for a heavy system of electrons and protons. As previously anticipated, the effective mass $m^{\rm eff}_{e}\gg m_{e}$ replaces $m_{e}$ in the dynamic equations (4.3) and (5.8). As a consequence, the large temperature $T_{\gamma}>10m_{e}$ radiation flux $F_{\gamma}$ acting on an electron is necessarily required to accelerate a heavy system of both electrons and protons.
Such a dynamical configuration should be realised in an opaque radiation fluid of photons and pairs at large temperatures $T_{\gamma}\gg m_{e}$ and number $N_{\gamma}\gg\bar{N}_{e}=\bar{N}_{p}$, where the bulk numbers $\bar{N}_{e}$ and $\bar{N}_{p}$ of fully ionised electrons and protons are small contaminations. Otherwise, if the radiation fluid temperature $T_{\gamma}$ is not large enough, instead of accelerating protons, the bulk electrons $\bar{N}_{e}$ are accelerated backwards to massive protons by the induced electric field $E_{pe}$, as numerically studied in Ref. [27].
In addition, when electrons and protons drift through the radiation fluid of photons and pairs, they undergo Klein-Nishina scattering processes. The KN cross section $\sigma^{p}_{{}_{\rm KN}}$ of drifting protons is much smaller than the one $\sigma^{e}_{{}_{\rm KN}}$ (4.1) of drifting electrons, i.e., $\sigma^{p}_{{}_{\rm KN}}\sim{\mathcal{O}}[(m_{p}/m_{e})^{2}]\sigma^{e}_{{}_{\rm KN}}$. Therefore, we neglect the cross section $\sigma^{p}_{{}_{\rm KN}}$ of drifting protons in the Klein-Nishina scattering processes between an electron-proton pair and the radiation fluid of photons and pairs. In these approximations, all analyses and results in Section 4 remain the same. Both drifting electrons and protons run away, becoming UHE particles. Figures 5 also represent the fraction of accelerated protons with the substitution $e\rightarrow p$. In the following, without further specification, all discussions and calculations apply to a bound system of electrons and protons with an effective electron mass $m^{\rm eff}_{e}=m_{p}+m_{e}$. Note that for the case $\gamma_{e}\approx\gamma_{p}$ the accelerated protons’ energy $\gamma_{p}m_{p}$ is ${\mathcal{O}}[(m_{p}/m_{e})]$ larger than the accelerated electrons’ energy $\gamma_{e}m_{e}$.
6 Luminosities of UHE emission and blackbody radiation
The accelerated electrons $N_{e}$ and protons $N_{p}$ reach ultra-high energies and stream out of the fireshell, carrying away the thermal energy $E_{\gamma}$ (2.6) of the radiation fluid of photons and pairs. We will calculate the luminosity of ultra-high-energy emissions of electrons and protons, as well as the time variation of the fireshell thermal energy $E_{\gamma}$. As will be calculated below, the luminosity and the rate $\dot{E}_{\gamma}/E_{\gamma}$ of thermal energy dissipation are small. Up to a certain timescale, the processes of thermal energy transfer and temperature cooling are nearly adiabatic, and the fireshell thermal equilibrium is maintained with slowly decreasing temperature. Therefore, we regard the fireshell as an approximate steady energy $E_{\gamma}$ reservoir, like the Sun. As a result of thermal energy dissipation and strong gravitational effect, the fireshell eventually cools down and shrinks in size.
6.1 Blackbody radiation luminosity and MeV annihilation line
Due to the large opacity and temperature gradient in the centre of the fireshell ($r_{+}<r<2r_{+}$), internal photons diffuse outwards in the radial direction. At the fireshell outer boundary $r=2r_{+}$, where the temperature is about $2m_{e}$ and pair annihilation occurs, it becomes optically thin, and photons stream out as a continuous blackbody radiation of the characteristic temperature $2m_{e}$. One obtains the blackbody luminosity by solving the radiation transfer (photon diffusion) equation [34],
$L^{0}_{\gamma}$ $\sim$ $4\pi(2r_{+})^{2}\sigma_{{}_{\rm SB}}d_{f}(T_{\gamma}^{+})^{4}c\left(\frac{\xi_{\gamma}}{2r_{+}}\right)\approx 4\left(\frac{\xi_{\gamma}}{2r_{+}}\right)L^{+}_{\gamma}$. (6.1)
The continuous radiation time scale is
$\frac{E_{\gamma}}{L^{0}_{\gamma}}$ $\approx$ $\left(\frac{1}{4\eta}\right)\left(\frac{\delta\ell}{c}\right)\left(\frac{2r_{+}}{\xi_{\gamma}}\right),$ (6.2)
where the photon luminosity $L^{+}_{\gamma}=\eta L_{\rm pl}$ (2.12) and mean-free path $\xi_{\gamma}\sim 10^{-4}\lambda_{e}$ (Fig. 3) are inside the centre of the fireshell of photons and pairs. In this case, the screen factor $\xi_{\gamma}/2r_{+}\sim 8.3\times 10^{-22}$ is small for the fireshell size $\sim 2r_{+}$ (2.4). The blackbody radiation luminosity (6.1) is a constant $L^{0}_{\gamma}\sim 10^{-22}L_{\rm pl}$, and the time scale (6.2) is large. These values depend on the fireshell parameters (2.4), (2.5), (2.10), and (2.12).
This emission is analogous to the blackbody radiation from the Sun’s surface. To compare and contrast, we mention the Sun’s surface halo of the temperature $\sim 6\times 10^{3}C^{\circ}\approx 0.55$ eV and the Sun radius $6.96\times 10^{10}$cm, where electrons are not relativistic ($\langle\gamma_{e}\rangle\approx 1$) and their densities are much smaller than the fireshell under discussion. The Sun’s luminosity is $L_{\odot}=3.83\times 10^{33}{\rm ergs/sec}\approx 10^{-27}L_{\rm pl}$. Thermal photons diffuse from the Sun’s core at the temperature $T_{\odot}=15\times 10^{6}C^{0}=1.5\times 10^{4}{\rm eV}=3\times 10^{-2}m_{e}$. While in the fireshell case, the blackbody has a characteristic temperature $\sim 2m_{e}$. Moreover, the radiation spectrum contains the electron-positron annihilation line and reheating of temperature increment $(11/4)^{1/3}$ before and after annihilation around $2m_{e}\approx 1$ MeV [14].
Figure 6: For the purpose of illustrating the spectral behaviour of UHE particle luminosity, we select the parameters: the horizon temperature (2.10) $T_{\gamma}^{+}=10m_{e},20m_{e}$ and acceleration number $A=10^{-3}$ (6.4), to plot high-energy electron and proton luminosities $L_{e}/L_{\rm pl}$ (6.3) in terms of the particle Lorentz factor $\gamma_{e}$.
6.2 UHE emission luminosity and energy spectrum
Section 4 shows that the KN cross section (4.1) decreases as accelerated electron and proton energies increase. Therefore, we expect high-energy electrons to stream out and carry away the fireshell energy $E_{\gamma}$ of photons and pairs. Using (i) the rates of radiation energy transfer to an accelerated electron and proton pair (5.10), (ii) total bulk electron number $\bar{N}_{e}=BN_{\gamma}$ (2.13) and (iii) photon-pair number $N_{\gamma}$ (2.11), we obtain the luminosity of high-energy electrons (protons) of the Lorentz factor $\gamma_{e}\approx\gamma_{p}$,
$L_{e}$ $=$ $\frac{dE_{e}}{dt}\approx N_{e}\sigma_{{}_{T}}F_{\gamma}(r_{+})[(4/3)(\beta_{e}\gamma_{e})^{2}]$ (6.3)
$=$ $A\left(\frac{N_{e}}{\bar{N}_{e}}\right)\left(\frac{n^{+}_{\gamma}}{m^{3}_{e}}\right)(4/3)(\beta_{e}\gamma_{e})^{2}L_{\gamma}^{+},$
where the “acceleration number” is defined as
$A=B(\sigma_{{}_{T}}\delta\ell m_{e}^{3}),$ (6.4)
indicating the effective number of bulk electrons $\bar{N}_{e}$ accelerated inside the volume $(\sigma_{{}_{T}}\delta\ell)$ of one-dimensional tube of photon-electron interactions. We treat the $A$ as a parameter, depending on the baryon matter contamination $B\ll 1$, the firespot size $\delta\ell$ and $(\sigma_{{}_{T}}\delta\ell m_{e}^{3})=(8\pi/3)\alpha^{2}(\delta\ell/\lambda_{e})$, which can be much larger than one.
We point out that the obtained UHE emission luminosity (6.3) exhibits three distinct features:
  1. (1)
    The luminosity sensitively depends on the horizon temperature $T_{\gamma}^{+}$ (2.10), via the photons and pairs number density $n^{+}_{\gamma}\equiv n_{\gamma}(T_{\gamma}^{+})$ (2.9) and the fraction $(N_{e}/\bar{N}_{e})$ (4.15) of electrons accelerated up to the Lorentz factor $\gamma_{e}$. Figure 5 shows that the fraction $N_{e}/\bar{N}_{e}$ is exponentially suppressed by small temperatures $T^{+}_{\gamma}\lesssim 10m_{e}$, therefore, the UHE emission luminosity (6.3) is completely negligible in these cases, unless the non-trivial acceleration number $A\gg 1$ (6.4), which relates to the baryon-loading $B$ and size $\delta\ell$ of the fireshells (spots) of energetic photons and pairs.
  2. (2)
    The fraction $N_{e}/\bar{N}_{e}$ properties (4.16), (4.17) and Figure 5 imply the UHE emission luminosity (6.3) of particles’ energies $m_{e}^{\rm eff}\gamma_{e}$ follows the spectral behaviour
    $(i)~L_{e}\propto(\gamma_{e})^{2-\nu},~(\gamma_{e}<10^{3});~~~(ii)~L_{e}~\propto(\gamma_{e})^{2},~(\gamma_{e}>10^{3}),$ (6.5)
    as shown in Fig. 6. This is consistent with the Compton-rocket effect (3.1) and (3.2) [21], showing that relativistic electrons gain more energy ($\gamma_{e}>1$) from radiation photons than non-relativistic ones ($\gamma_{e}\approx 1$).
  3. (3)
    Similar to the blackbody radiation (6.1), the UHE emission is not a transient phenomenon, but rather a lasting, continuous radiation. The reason is that the fraction $(N_{e}/\bar{N}_{e})$ becomes tiny, and the emission time scale
    $\frac{E_{\gamma}}{L_{e}}$ $\approx$ $\left(\frac{\delta\ell}{c}\right)\left[A\left(\frac{N_{e}}{\bar{N}_{e}}\right)\left(\frac{n_{\gamma}^{+}}{m^{3}_{e}}\right)(4/3)(\beta_{e}\gamma_{e})^{2}\eta\right]^{-1},$ (6.6)
    becomes even larger than the blackbody one (6.2).
In the second point (2), the luminosity increases with the UHE particles’ energy $m_{e}^{\rm eff}\gamma_{e}$, but the increment in luminosity $L_{e}\propto(\gamma_{e})^{2}$ with $\gamma_{e}$ does not run to infinity $L_{e}\propto(\gamma_{e})^{2}\rightarrow\infty$ in time. The reason is due to the back reaction that the large luminosity $L_{e}$ of large $\gamma_{e}$ decreases the fireshell temperature $T_{\gamma}^{+}$ and energy $E_{\gamma}$ fast in time, as required by total energy conservation. As a result, the accelerated electron fraction $N_{e}/\bar{N}_{e}$ (4.18) and the UHE emission luminosity $L_{e}$ (6.3) of large $\gamma_{e}$ decrease fast in time, as will be shown below.
To compare two emission processes from the fireshell, we calculate the ratio of the blackbody radiation (6.1) and UHE emission (6.3) luminosities given by
$\frac{L_{e}}{L^{0}_{\gamma}}$ $=$ $A\left(\frac{N_{e}}{\bar{N}_{e}}\right)\left(\frac{n^{+}_{\gamma}}{m^{3}_{e}}\right)(\beta_{e}\gamma_{e})^{2}\left(\frac{\xi_{\gamma}}{2r_{+}}\right)^{-1}.$ (6.7)
There are two possibilities:
  1. (i)
    the blackbody radiation is dominant $L_{e}/L^{0}_{\gamma}\ll 1$ for low temperature $T^{+}_{\gamma}\gtrsim m_{e}$ and small fraction $N_{e}/\bar{N}_{e}\ll 1$;
  2. (ii)
    the UHE emission is dominant $L_{e}/L^{0}_{\gamma}\gg 1$ for large temperature $T^{+}_{\gamma}\gg m_{e}$ and the significantly non-trivial fraction $N_{e}/\bar{N}_{e}$.
The former case (i) is well studied, and the blackbody luminosity is small due to the fireshell’s large opacity $(r_{+}/\xi_{\gamma})\gg 1$. We are primarily interested in the latter case (ii) for $\gamma_{e}\gg 1$, when $T^{+}_{\gamma}\gg m_{e}$ and the UHE emission dominantly dissipates the energy $E_{\gamma}$ (2.6) of the fireshell.
Figure 7: For illustrating the numerical results, we adopt the initial horizon temperature $T_{\gamma}^{+}(0)=100m_{e}\approx 50$ MeV (2.10) and the acceleration number $A=10^{-3}$ (6.4). We consider the Lorentz factor $\gamma_{e}\sim 10^{9}$ channel, corresponding to UHE electron energies $\sim 10^{3}$ TeV and proton energies $\sim 10^{3}$ PeV. The Compton time $\tau_{e}=(\lambda_{e}/c)=1.28\times 10^{-21}{\rm sec}$, and the final time $10^{26}\tau_{e}$ is about days. The $L_{e}/L_{\gamma}^{0}$ broken line (log-log plot) indicates the transition tuning place from the UEH emission dominant $L_{e}/L^{0}_{\gamma}\gg 1$ to the blackbody ($B$) emission dominant $L_{e}/L^{0}_{\gamma}\ll 1$. For different values of $\gamma_{e}$ and $T_{\gamma}^{+}(0)$, the results will exhibit the same behaviours in time variations. The UHE emission luminosity $L_{e}$ increases for large initial temperature $T_{\gamma}^{+}(0)$ and acceleration number $A$. The emission duration lasts longer for smaller $\gamma_{e}$ channels of UHE particles.
6.3 Fireshell temperature and UHE emission decrease in time
The blackbody emission (6.1) and the UHE emission (6.3) dissipate the fireshell thermal energy $E_{\gamma}$ (2.6), therefore, the total energy conservation yields
$-\frac{dE_{\gamma}}{dt}=\frac{dE^{0}_{\gamma}}{dt}+\frac{dE_{e}}{dt}=L^{0}_{\gamma}+L_{e}.$ (6.8)
We use this equation together with Eqs. (2.6), (6.3) and (6.7) to determine how the horizon temperature $T^{+}_{\gamma}(t)$ decreases in time,
$\frac{dT^{+}_{\gamma}(t)}{dt}$ $=$ $-1.23B\left(\frac{N_{e}}{\bar{N}_{e}}\right)\left(\frac{\sigma_{{}_{T}}}{4\pi r_{+}^{2}}\right)(\beta_{e}\gamma_{e})^{2}L_{\gamma}^{+}\left(1+\frac{L_{\gamma}^{0}}{L_{e}}\right)$ (6.9)
$=$ $-10.4\eta A\left(\frac{N_{e}}{\bar{N}_{e}}\right)(\beta_{e}\gamma_{e})^{2}m_{e}^{2}\left(1+\frac{L_{\gamma}^{0}}{L_{e}}\right),$
where the factor $\eta$ in the second line comes from the boundary condition (2.4). In the fireshell decay equation (6.9), we particularly consider a large $\gamma_{e}\gg 1$ channel, because it dominates the energy dissipation of the fireshell. Before presenting numerical calculations, we note that the gravitational redshift factor in energy and the gravitational dilation in time are simplified in the luminosities $L_{\gamma}^{0}$ (6.1) and $L_{e}$ (6.3), and the temperature variation $dT^{+}_{\gamma}/dt$ (6.9).
For a given channel $\gamma_{e}\gg 1$ value and an initial value $T^{+}_{\gamma}(0)$ at the time $t=0$, we use the numerical result $N_{e}/\bar{N}_{e}$ (4.15) as a function of $T^{+}_{\gamma}(t)$ to integrate the fireshell decay equation (6.9) over the time. Furthermore, substituting the numerical solution $T^{+}_{\gamma}(t)/m_{e}$ into $(N_{e}/\bar{N}_{e})$ and $(n_{\gamma}^{+}/m_{e}^{3})$ in Eq. (6.3), we calculate the UHE emission luminosity $L_{e}(t)$ declining as a function of the time $t/\tau_{e}$. We calculate these integrations numerically by using the Mathematica program. Figure 7 shows that the numerical results of the horizon temperature $T^{+}_{\gamma}(t)/m_{e}$, the accelerated electron fraction $N_{e}(t)/\bar{N}_{e}$, and the UHE emission luminosity $L_{e}(t)$ monotonically decrease in time $t/\tau_{e}$ varying up to $10^{26}\tau_{e}$, which is about days. The results show two phases:
  1. (i)
    The initial phase: the UHE emission initially dominates over the blackbody radiation $L_{e}/L_{\gamma}^{0}>1$, and the turning point to $L_{e}/L_{\gamma}^{0}<1$ occurs around $t=10^{25}\tau_{e}$ when the temperature $T^{+}_{\gamma}\approx 10m_{e}$, and the accelerated electron fraction $N_{e}(t)/\bar{N}_{e}$ decreases significantly in time.
  2. (ii)
    The latter phase: the fireshell cooling process attributed to the blackbody radiation $L_{\gamma}^{0}$ (6.1) and the temperature $T^{+}_{\gamma}(t)$ decreases in time at a different slope (power law) from the one in the initial phase.
In the UHE emission luminosity-dominant phase $L_{e}/L^{0}_{\gamma}\gg 1$, we neglect the small time-variation effect of the blackbody luminosity $L_{\gamma}^{0}(t)$ (6.1) caused by the time varying mean-free path $\xi_{\gamma}=(\sigma_{\gamma}n^{+}_{\gamma})^{-1}$ (2.8) and number density $n^{+}_{\gamma}\propto[T^{+}_{\gamma}(t)/m_{e}]^{3}$ of photons and pairs in the fireshell. The effect should probably be considered in the blackbody luminosity $L_{e}/L^{0}_{\gamma}\ll 1$ dominant phase, when the horizon temperature $T^{+}_{\gamma}<10m_{e}$, electron-positron pair annihilation into two photons occurs, and the fireshell opacity decreases. As a consequence, the TOV equilibrium (2.3) is gradually lost and the fireshell (or fire spot) shrinks and disappears. These are not the topics of this study and will be subjects of future investigations.
7 Characteristics relevant for observations and experiments
7.1 Peculiar features of UHE particles produced in such a scenario
We highlight some qualitative characteristics of UHE particles produced by the scenario discussed in previous sections:
  1. (1)
    The mechanism occurs at local microscopic length and time scales, over which the photon-pair fluid temperature ($T_{\gamma}\gg m_{e}$) has a large gradient, resulting in a strong radiation force for electron and proton acceleration. Therefore, the mechanism does not essentially depend on the global symmetry, macroscopic length and time scales of high-energy astrophysical environments and strong gravitational processes, which create such a fire spot of highly dense and energetic fluid of photons and pairs. These fire spots contain thermal energy $E_{\gamma}$ (2.6) of the radiation fluid of photons and pairs, which are energetic sources for UHE particle productions. Therefore, the trapped fireshell (halo) near the horizon is only an example to simplify calculations and illustrations. The fire spots can be formed, for instance, in gravitational binary coalescence and accretion cases, in the early Universe, and may be observed in future laser experiments.
  2. (2)
    The horizon temperature $T^{+}_{\gamma}>10m_{e}$, as if an effective overcritical classical An electric field $E_{\rm eff}\sim{\mathcal{O}}(10)E_{c}$ provides a significant probability of accelerating electrons and protons to UHE, up to TeV, PeV, EeV, and even higher. The total luminosity of UHE particles increases as the acceleration number $A$ increases, namely, more electrons in the fireshell participate in the acceleration process. The UHE particle luminosity (6.3) could be significant for small temperatures $2m_{e}<T^{+}_{\gamma}<10m_{e}$ and large acceleration numbers $A$. However, we cannot show the results, because the numerical precision is limited in calculating the exponential factor $e^{-m_{\rm eff}/T_{\gamma}^{+}}$.
  3. (3)
    The UHE emission is not a transient phenomenon, such as a burst, but a long-lasting continuous radiation. UHE particles have peculiar spectral behaviours in high and low energies: approximate power laws $\propto\gamma^{-\nu}_{e}$ for small Lorentz factor $\gamma_{e}\lesssim 10^{3}$, and $\propto\gamma_{e}^{2}$ for larger $\gamma_{e}$. The luminosity of UHE particles should be suppressed for very large $\gamma_{e}$ due to the back reaction of UHE particle production on the sources of fire spots.
  4. (4)
    Two fire-spot parameters are essential for characterising different sources of producing UHE particles:
    $(T^{+}_{\gamma},~A).$ (7.1)
    The temperatures $T^{+}_{\gamma}$ of fire spots of photons and pairs are crucial for UHE particle spectra ($\gamma_{e}$). The acceleration number $A$ of electrons (protons) participating in acceleration dynamics is relevant to the total luminosity of UHE particles. For the appropriate values $T^{+}_{\gamma}$ and $A$ of the fire spots of astrophysical sources, the resultant UHE emission luminosity (emission probability) and energy spectra (Lorentz factor $\gamma_{e}$ or $\gamma_{p}$) can be relevant for the observations of UHE particles. In general, astrophysical fire-spot sources with large (small) $T^{+}_{\gamma}$ and $A$ values produce large (small) luminosity and energy of UHE particles. While the sources with large $T^{+}_{\gamma}$ and $A$ values are less probable. This may explain why the observed UHE particle luminosity monotonically decreases with energy.
In principle, the primary UHE electrons and protons are energetic enough to produce secondary VHE neutral photons and neutrinos through their decays and/or interactions with low-energy particles in the environments of sources of fire spots. Namely, UHE electrons and protons interact with low-energy fields and particles in the medium around their sources, producing VHE photons and neutrinos. Since these interactions are highly energetically forward scattering, resultant secondary VHE particles carry the characteristics of primary UHE particles. This implies that UHE and VHE particle events should be observationally related to each other and correlate with astrophysical sources.
7.2 Possible relevance for observations and experiments
In this article, we focus only on revealing the novel scenario and mechanism of generating UHE particles, and pay no attention to quantitatively discussing this scenario compared with the observations of UHE and VHE particles. Nevertheless, we mention some observationally and experimentally relevant features, in particular for UHE and VHE particles accompanying GRB phenomena observed.
Interacting with magnetic fields, soft photons and other low-energy particles around sources, these UHE electrons and protons from the opaque radiation fluid spot of highly energetic photons and pairs produce VHE photons and neutrinos to TeV or PeV energy scales. We come to GRBs’ cases. As shown in Fig. 7, the trapped fireballs (horizon halos) cool via continuous emission of UHE electrons and protons. Interacting with the surrounding fields and materials, the former mainly leads to VHE photons, and the latter to VHE neutrinos. This scenario strongly implies the logical and causal reasoning features that are relevant to the observations of VHE photons accompanying GRBs’ phenomena:
  1. (1)
    UHE electrons (protons) and the subsequent VHE photons and neutrinos are continuous emissions in the fireshell lifetime. They should correlate with each other with the GRBs’ sources in observations.
  2. (2)
    Before outward expanding fireballs become optically thin, we probably cannot see VHE photons because of great opacity due to two-photon annihilation $\gamma\gamma\rightarrow e^{+}e^{-}$ of the Breit-Wheeler process (2.7). During the prompt emissions for bursts, the VHE photons most probably observed are after expanding fireballs become transparent, namely, VHE photon luminosity increases soon after the main GRBs’ bursts occur  [^8] The features (1) and (2) were pointed out in Ref. [17].
  3. (3)
    During the afterglow era, UHE protons $\gamma_{p}\gg 1$ emitted from the horizon fireshell continuously collide with the ejector of the Lorentz factor $\Gamma<\gamma_{p}$, which then emits VHE photons and neutrinos. Therefore, their (VHE) light curves follow the law of the ejector’s kinematic motion in time, similarly to the external shock emission of the ejector colliding with the interstellar medium. However, their (VHE) energy spectra become softer in time because the horizon fireshell cool down. The time and luminosity of UHE and VHE continuous emissions depend on the horizon fireshell temperature $T_{\gamma}$ and acceleration number $A$ (7.1), as well as the target (ejector) density.
In contrast, according to the usual modelling interpretation of the TeV emission in GRBs, ultra-relativistic electrons are accelerated by internal shocks of fireballs during the prompt emission or by external shocks between the ejector and interstellar medium (ISM) during the afterglow phase.
In addition to astrophysical sources, the scenario and mechanism in principle apply to the early Universe at large particle density and high temperature $T_{\gamma}>2m_{e}$ before the Universe becomes transparent. Local dense and inhomogeneous radiation fields accelerate charged particles to high energies. It implies that in the cosmic background observed today, there should be imprints of excess UHE and VHE particles originating from the early Universe. Moreover, such an energetic spot would be created in future by strong laser-beam experiments in ground laboratories. The probability of producing UHE particles could be non-trivial for detection, if the strong laser field spot is not only sufficiently energetic $T_{\gamma}>10m_{e}$ and dense $A\gg 1$, but also has a lifetime longer than the time scale $\tau^{e}_{{}_{\rm KN}}=(\sigma^{e}_{{}_{\rm KN}}n_{\gamma}c)^{-1}$ (4.2) for the runaway process.
8 Summary and remarks
In this article, we study the possibility of how electrons and protons are accelerated to UHE particles in an opaque fire spot of dense photons and $e^{-}e^{+}$ pairs at a large temperature $T_{\gamma}\gg m_{e}$. We have revealed a possible acceleration mechanism for UHE charged particles, which mainly attribute to (i) the Compton-rocket effect of strong radiation force acting on electrons for acceleration; (ii) the instability and runaway dynamics occur when accelerated electrons undergo the KN scattering process, whose cross section decreases with increasing electron energy; (iii) such drifting electrons develop a local electric field acting on protons for acceleration. For qualitative analysis and theoretical illustrations, we use the Compton-rocket equation for electrons’ energy gain and the rate equation of electrons’ energy loss in Thomson and KN scatterings in a simplified one-dimensional model. As a result, we obtain the non-trivial probability of accelerating electrons and protons reaching a large Lorentz factor $\gamma_{e}\sim 10^{9}$ over a small distance $d\sim 10^{-2}$ cm.
As an astrophysical source for the opaque fire spots of dense photons and pairs at large temperatures, we consider the opaque fireballs formed in gravitational collapses. The outer parts of fireballs expand, accounting for the phenomena of GRBs. Whereas the inner parts of fireballs, fireshell, slowly contract and get dense near the horizon due to strong gravity and large pressure. The inner part of fireballs, fireshell as an energy source, possesses large opacity and thermal energy with a small amount $B$ of electrons and protons from baryon matter contamination, which are comparable with counterparts of the outer part of fireballs. Approximating the fireshell as a metastable state, we calculate the luminosities and spectra of UHE emission and blackbody radiation. In this scenario, we point out distinct features of produced UHE and VHE particles, which are relevant for observations and in contrast with the usual interpretations of the TeV emission in GRBs.
However, we emphasise that the systems and dynamics under investigation are rather complex, consisting of photons and pairs, protons and electrons, various interacting processes, whose times and lengths run from the microscopic to the macroscopic scales. In these preliminary studies, the models and conditions used to present the idea and scenario are simplified, the fraction $N_{e}/\bar{N}_{e}$ and the analyses used to show the results are qualitative. The one-dimensional electric field between drifting protons and electrons should be generalised to the three-dimensional case. Full interaction rates and cross sections depending on energy and momentum should be considered to investigate how the instability of electron perturbations develops in the strong radiation force of dense and energetic photons and pairs, leading to an avalanche runaway process. Needless to say, it is necessary to develop a numerical code for performing full simulations to verify the present scenario of accelerating electrons and protons to ultra-high energies. The dynamics are attributed to the radiation force in an opaque fire spot of dense photons and pairs at temperatures that exceed the electron mass and have a large spatial gradient in a preferential direction. Moreover, it is also necessary to study how such firespots can be established in high-energy astrophysical processes in strong gravitational environments, and determine their temperatures, sizes and lifetimes. In observations, it is worthwhile to phenomenologically examine the three distinct features of VHE photon emissions in GRBs mentioned in Sec. 7.2, and possible relations between observed cosmic ray events of UHE particles and VHE emissions of photons and neutrinos, and correlations with high-energy astrophysical sources.